Question 2
The Blair Company’s three assembly plants are located in California, Georgia and New
Jersey. Previously, the company purchased a major subassembly, which becomes part
of the final product, from an outside firm. Blair has decided to manufacture the
subassemblies within the company and must now consider whether to rent one centrally
located facility (e.g., in Missouri, where all subassemblies would be manufactured) or to
rent three separate facilities, each located near one of the assembly plants, where each
facility would manufacture only the subassemblies needed for the nearby assembly
plant. A single, centrally located facility, with a production capacity of 18,000 units per
year, would have fixed costs of $900,000 per year and a variable cost of $250 per unit.
Three separate decentralized facilities, with production capacities of 8,000, 6,000 and
4,000 units per year, would have fixed costs of $475,000, $425,000 and $400,000,
respectively, and variable costs per unit of only $225 per
1
Expert's answer
2013-02-18T10:11:43-0500
For the single plant: TC = FC + VC = 900,000 +250* (6,000 + 4,500 + 3,000) = 900,000 + 3,375,000 = $4,275,000
For 3 plants: TC = FC + VC = (475,000 + 425,000 + 400,000) + 225*(6,000 + 4,500
+ 3,000) = 1,300,000 + 3,037,500 = $4,337,500
So, for current level of production a single plant is better. But if the level
increase, the current way of production is better.